# OSS License Identification at Scale: A Comprehensive Dataset Using World of Code thanks: Replication package available at: https://zenodo.org/records/14279932

### v2

@@ Abstract
The proliferation of open source software (OSS) and different types of reuse has made it incredibly difficult to perform an essential legal and compliance task of accurate license identification within the software supply chain. This study presents a reusable and comprehensive dataset of OSS licenses, created using the World of Code (WoC) infrastructure. By scanning all files containing “license” in their file paths, and applying the approximate matching via winnowing algorithm to identify the most similar license from the SPDX list, we found and identified 5.5 million distinct license blobs in OSS projects. The dataset includes a detailed project-to-license (P2L) map with commit timestamps, enabling dynamic analysis of license adoption and changes over time. To verify the accuracy of the dataset we use stratified sampling and manual review, achieving a final accuracy of 92.08%, with precision of 87.14%, recall of 95.45%, and an F1 score of 91.11%. This dataset is intended to support a range of research and practical tasks, including the detection of license noncompliance, the investigations of license changes, study of licensing trends, and the development of compliance tools. The dataset is open, providing a valuable resource for developers, researchers, and legal professionals in the OSS community.
@@ III-B License Blob Identification
The second category was acceptable, as we don’t claim a blob contains only the matched license. However, the third, with additional clauses, was concerning as it could alter the license’s nature. Detailed results are in Table II.
@@ III-C Project to License Mapping
To create the project-to-license (P2L) map, we use the 5.5 million matched license blobs and join them with WoC’s blob-to-time project (b2tP) map, which links blobs to the projects they were committed to, along with commit timestamps. This produces a table mapping each project to a known license and the time of the commit.
However, a blob’s presence in a project’s latest status can’t be confirmed solely from commit history, as it might have been removed later. To address this, we use the project-to-last-commit (P2lc) and tree-to-objects (t2all) maps from WoC. The P2lc map links projects to their last commit at the time of the latest WoC update (Version V), allowing us to retrieve the list of all blobs in a project’s current state by joining P2lc, c2dat (commit-to-tree), and t2all maps. This method not only provides all the times at which a blob was committed to a project but also verifies whether it still exists in the project.
@@ III-E Complementing Data
Although our P2L map already demonstrated strong performance in manual verification, we incorporated the previously published dataset by Gonzalez-Barahona et al. [4] to enhance data comprehensiveness. Their dataset includes only blobs and their detected licenses using ScanCode [14]. We filtered data to blobs with license detection confidence 95% or higher and applied the same process described earlier to map these blobs to commits and projects, enabling us to determine the time and project in which each license was committed. The merged table includes a column indicating the license detection method for each entry: either our method (1-WoC) or the Software Heritage dataset method (2-SH) [4].

### v3

@@ Abstract
The proliferation of open source software (OSS) and different types of reuse has made it incredibly difficult to perform an essential legal and compliance task of accurate license identification within the software supply chain. This study presents a reusable and comprehensive dataset of OSS licenses, created using the World of Code (WoC) infrastructure. By scanning all files containing “license” in their file paths, and applying the approximate matching via winnowing algorithm to identify the most similar license from the SPDX and Open Source list, we found and identified 5.5 million distinct license blobs in OSS projects. The dataset includes a detailed project-to-license (P2L) map with commit timestamps, enabling dynamic analysis of license adoption and changes over time. To verify the accuracy of the dataset we use stratified sampling and manual review, achieving a final accuracy of 92.08%, with precision of 87.14%, recall of 95.45%, and an F1 score of 91.11%. This dataset is intended to support a range of research and practical tasks, including the detection of license noncompliance, the investigations of license changes, study of licensing trends, and the development of compliance tools. The dataset is open, providing a valuable resource for developers, researchers, and legal professionals in the OSS community.
@@ III-B License Blob Identification
The second category was acceptable, as we do not claim a blob contains only the matched license. However, the third, with additional clauses, was concerning as it could alter the license’s nature. Detailed results are in Table II.
@@ III-C Project to License Mapping
To create the project-to-license (P2L) map, we use the 5.5 million matched license blobs and join them with WoC’s blob-to-time project (b2tP) map, which links blobs to the projects they were committed to, along with commit timestamps. This produces a table mapping each project to a known license and the time of the commit (see Figure 1).
However, a blob’s presence in a project’s latest status cannot be confirmed solely from commit history, as it might have been removed later. To address this, we use the project-to-last-commit (P2lc) and tree-to-objects (t2all) maps from WoC. The P2lc map links projects to their last commit at the time of the latest WoC update (Version V), allowing us to retrieve the list of all blobs in a project’s current state by joining P2lc, c2dat (commit-to-tree), and t2all maps. This method not only provides all the times at which a blob was committed to a project but also verifies whether it still exists in the project.
@@
Figure 1: License Identification Data Flow Diagram
@@ III-E Complementing Data
Although our P2L map already demonstrated strong performance in manual verification, we incorporated the previously published dataset by Gonzalez-Barahona et al. [4] to enhance data comprehensiveness. Their dataset includes only blobs and their detected licenses using ScanCode [14]. We filtered data to blobs with license detection confidence 95% or higher and applied the same process described earlier to map these blobs to commits and projects, enabling us to determine the time and project in which each license was committed. The merged table (see Figure 1) includes a column indicating the license detection method for each entry: either our method (1-WoC) or the Software Heritage dataset method (2-SH) [4].
@@ V Limitations
Acknowledgments
This work was supported in part by the National Science Foundation under Award Numbers 1901102 and 2120429.